OpenStudy (anonymous):

give exact and the approx. solutons to three decimal places x^2+7x-5=0. I am really mixing myself up!

OpenStudy (anonymous):

\[x^2+7 x-5=0 \]\[x^2+7 x=5 \]\[x^2+7 x+\left(\frac{7}{2}\right)^2=5+\left(\frac{7}{2}\right)^2 \]\[\frac{1}{4} (2 x+7)^2=\frac{69}{4} \]\[(2 x+7)^2=69 \]\[2 x+7=\pm \sqrt{69} \]\[2 x=-7\pm \sqrt{69} \]\[x=\frac{-7\pm \sqrt{69}}{2} \]

OpenStudy (anonymous):

Thank you! How do I get the approx. solution?

OpenStudy (anonymous):

\[x=0.5 (-7.\pm 8.30662)\]\[\text{Round}[\{0.5 (-7.+8.30662),0.5 (-7.-8.30662)\},0.001] \]\[\{0.653,-7.653\} \]

OpenStudy (anonymous):

ohh! I was busting my brains and I should've known that! Thanks for your help!!:)

OpenStudy (anonymous):

You are welcome.